If Sam can do the job in 6 days.. carla can do it in 4 days... John can do it in 2 days. How many days do they need .. if they do it together..
Found 2 solutions by scott8148, oberobic:
Answer by scott8148(6628)

they each do a fraction of the task based on their individual rates; the fractions sum to one, the whole task

(x/6) + (x/4) + (x/2) = 1

2x + 3x + 6x = 12

Answer by oberobic(2304)

The problems hinge on the fraction of work each can do per day, times the number of days.
.
(1/6 + 1/4 + 1/2)x = 1 whole job
.
(2/12 + 3/12 + 6/12)x = 1
.
11/12*x =1
.
11x = 12
.
x = 12/11
.
So working together they can do the job in 12/11 of a day.
.
Done.